# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v1

@@ 3.4.1. Problem Formalization
Given the test suite before minimization $\mathcal{T}$, the minimization problem is defined as a fixed-size subset selection with $100\%$ requirement coverage as the constraint. Each solution (i.e., subset $\mathcal{T}^{\prime}$) is represented as a binary vector:
@@
Given the minimization budget $\mathcal{M}\in(0,1]$, the number of test cases in the subset $\mathcal{T}^{\prime}$ is constrained by:
(5) $\sum_{i=1}^{m}x_{i}=\mathrm{round}\left(m\cdot\mathcal{M}\right)$
where $\mathrm{round}()$ denotes rounding the number to the nearest integer.
@@
Given that the objective of the search process is to minimize the similarity values between the pairs of selected test cases, the fitness function is formulated as:
(7) $Fitness=\frac{\sum_{i,t_{i}\in\mathcal{T}^{\prime}_{n}}(Max_{i,j,t_{i},t_{j}\in\mathcal{T}^{\prime}_{n},i\neq j}\ Norm\_Sim(t_{i},t_{j}))^{2}}{n}$
where $\mathcal{T}^{\prime}_{n}$ denotes the subset containing $n$ test cases, and $Norm\_Sim(t_{i},t_{j})$ represents the normalized similarity score between test cases pair $(t_{i},t_{j})$.
@@ 3.4.2. Initialization strategies
Our minimization problem is defined with two hard constraints: (1) the percentage of selected test cases in the subset should satisfy the minimization budget, and (2) the subset should preserve $100\%$ requirement coverage. In order to efficiently guide the search to find valid subsets, we devise three different strategies to initialize a set of valid subsets for the GA to use as a base for further search. Note that in our dataset, the relationship between test cases and requirements is many-to-one. Consequently, one limitation of the minimization problem with full requirement coverage is that the pre-defined number of selected test cases must not be smaller than the total number of requirements.
  • Strategy 1. Iterative Selection This strategy iteratively selects one test case randomly under each requirement, continuing this process until the total number of selected test cases reaches the minimization budget.
  • Strategy 2. Initial Requirement + Random
@@
    • Step 2: Randomly select the remaining test cases without considering requirements until the total number of selected test cases reaches the minimization budget.
  • Strategy 3. Proportional Selection
@@
    • Step 2: If the total number of selected test cases across all requirements is less than or greater than the minimization budget, iteratively add or reduce the number of selected test cases per requirement, ensuring every requirement is still covered, until the desired minimization budget is reached.
3.4.3. Evolutionary Search
After initializing a population of valid subsets using the strategies described in Section 3.4.2, we apply a GA to iteratively evolve the population toward solutions that minimize redundancy while preserving full requirement coverage and complying with the minimization budget.
Each individual in the population is encoded as a binary vector, as defined in Equation 4. The fitness of an individual is computed using the objective function described in Section 3.4.1.
At each generation, we perform the following operations:
@@ 3.4.3. Evolutionary Search
  • Crossover: We used the customized crossover operator (Blank and Deb, 2020) that ensures the offspring maintains a fixed subset size. It first includes all test cases shared by both parent subsets, then adds the test cases present in either parent subset until the minimization budget is met.
  • Mutation: We used an inversion mutation operator (Blank and Deb, 2020) that randomly selects a segment of the subset and reverses its order. Since each subset is represented as a binary vector, this approach preserves the subset size.
@@
We follow the published guidelines  [^1] https://www.obitko.com/tutorials/genetic-algorithms/recommendations.php used in prior work (Pan et al., 2024) for setting GA hyperparameters. Specifically, we use a population size of $100$, a mutation rate of $0.01$, and a crossover rate of $0.90$. For each generation, the size of each solution (i.e., subset) remains fixed and equal to the minimization budget. The evolutionary process is repeated until the convergence, which is defined as the improvement in the fitness score being less than $0.0025$ across generations. The final output is the best subset identified during the search, which satisfies all constraints and minimizes the redundancy among selected test cases, indicated by the lowest fitness value.
@@ 1. RQ1:
    This RQ aims to identify the optimal configuration of RTM that achieves the best performance (i.e., FDR). Specifically, a RTM configuration entails the selection of a preprocessing method (Section 3.1), an embedding technique (Section 3.2), a similarity measure (Section 3.3), and an initialization strategy (Section 3.4.2).
@@ 2. RQ2:
    How does RTM compare to state-of-the-art TSM techniques?
    This RQ aims to compare the performance of RTM with two baseline approaches, namely Random Minimization and FAST-R (Cruciani et al., 2019), under various minimization budgets (i.e., 10%, 20%, …, 90%).
@@ 4.2. Dataset
We evaluated RTM and compared it to baselines on a dataset collected from our industry partner, an automotive system company. There are $736$ test cases across seven test runs. As shown in Figure 1, the test cases consist of test steps that validate the behavior of the automotive system related to a specific DTC. Each test case covers one requirement, with a total of $54$ requirements covered, and may detect multiple faults, with a total of $220$ unique faults detected.
Note that we identified the faults, which are the root causes of the failing test cases (e.g., the parameter values mismatched, the connection error of specific components), using the test execution logs under the guidance of the system engineers. For each test case, we first extracted the failure messages from the test execution logs and then grouped the ones that were triggered by the same root cause into faults.
Test suites for different redundancy levels. In this dataset, one fault can be detected by multiple test cases, indicating a degree of redundancy in the test suite. We define the redundancy level for the test suite as follows:
@@
For example, for a test suite $\mathcal{T}$ with three test cases, each of which detects $3$, $5$, and $3$ faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is $4$, and the redundancy level of $\mathcal{T}$ is therefore $(3+5+3)/4=2.75$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
In order to assess the impact of the test suite redundancy level on the performance of RTM, we generate $10$ diverse test suites for each redundancy level using Integer Linear Programming (ILP)  [^2] https://pypi.org/project/PuLP/ and GA. This process starts by utilizing ILP to search for $100$ test suites with different test suite sizes while satisfying three constraints: (1) $100\%$ requirement coverage, (2) $100\%$ fault coverage, and (3) a specific test redundancy level. We set $15$ different redundancy levels, ranging from $4.5$ to $11.5$, with an interval of $0.5$. Note that there is no test suite that satisfies a redundancy level below $4.5$ under constraints (1) and (2), and the highest redundancy level is $11.86$, which is observed in the original test suite. Then, to enhance the diversity of the dataset for experiments, for the $100$ test suites under each redundancy level, we employ GA to search for the most diverse $10$ test suites that minimize the sum of the set similarity between test suites, which is defined as the Jaccard Similarity between two test suites:
@@
where $|T_{i}$cap $T_{j}|$ denotes the number of test cases in the intersection of test suite $T_{i}$ and $T_{j}$. $|T_{i}\cup T_{j}|$ denotes the number of test cases in the union of test suites $T_{i}$ and $T_{j}$.
@@ 4.3. Text Representations
Inspired by prior work on vector representations of requirements (Zhao and Alhoshan, 2025; Abbas et al., 2025) and test cases (Viggiato et al., 2022), we employ a variety of text embedding techniques to convert test cases into vector representations. Specifically, we use TF-IDF, USE, Amazon Titan Text Embedding V2, and LongT5 to generate sentence-level embeddings, while Word2Vec, GloVe, and FastText are used to obtain word-level embeddings. Due to limitations in computational resources, data privacy constraints, and input token length restrictions, we were unable to experiment with models such as BERT (Devlin et al., 2019) and Sentence-BERT (Reimers and Gurevych, 2019).
@@
TF-IDF. TF-IDF has demonstrated promising performance in identifying similar test cases expressed in natural language (Viggiato et al., 2022). We also employed TF-IDF to extract numerical vector representations of the test cases. For each word, we calculated its importance within a single test case relative to its occurrence across all other test cases.
@@
USE (Universal Sentence Encoder) (Cer et al., 2018). USE encodes text into high-dimensional vectors, capturing semantic information at the sentence level, making it particularly effective for tasks such as semantic similarity, text classification, clustering, and information retrieval (Cer et al., 2018).
@@
LongT5 (Guo et al., 2022) LongT5 is an extension of the T5 architecture designed to efficiently handle long sequences (Guo et al., 2022). It incorporates sparse attention and reversible layers to reduce memory usage, making it well-suited for tasks like long-document summarization and retrieval-augmented generation. We selected the local attention mechanism for LongT5 as it yielded a higher FDR for TSM compared to the global attention mechanism.
@@
Amazon Titan Text Embedding V2 (AWS, 2024). For LLM-based embedding, we use Amazon Titan Text Embedding V2, which is provided by our industrial partners. The model takes test case steps as input and produces a $1,024$-dimensional vector representation for test cases.
@@
Word2Vec (Mikolov et al., 2013). Word2Vec has been shown to be effective for identifying similar test cases (Viggiato et al., 2022). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ as the configuration for Word2Vec, as this setup yielded the highest FDR result for TSM. This was determined with experiments in which we evaluated both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
@@
GloVe (Pennington et al., 2014). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (Pennington et al., 2014). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of $2$ for GloVe according to its performance in the hyperparameter tuning experiments conducted with window sizes ranging from $2$ to $10$ in increments of $2$. Same as Word2Vec, we used GloVe to generate $300$-dimensional numerical vector representation for each word in the test cases.
@@
FastText (Bojanowski et al., 2017). FastText is a word embedding model proposed by Bojanowski et al. (Bojanowski et al., 2017), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of $8$ as the configuration for FastText, according to the results of experiments in which we evaluated both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
@@ 4.4. Baselines
We compare RTM to Random Minimization and FAST-R (Cruciani et al., 2019).
@@ 4.4.1. Random Minimization
Since the goal of TSM is to reduce the number of test cases and, consequently, lower testing costs, the most straightforward approach is to randomly remove test cases to meet the minimization budget (Ali et al., 2010). We compare the performance of RTM against random minimization, with and without considering the requirement coverage constraint. For the version with the constraint, we applied the same selection strategy as GA initialization described in Section 3.4.1. For the version without the constraint, we randomly select $k$ test cases from the test suite while ensuring $k$ adheres to the minimization budget.
4.4.2. FAST-R (Cruciani et al., 2019)
FAST-R is a family of similarity-based TSM approaches, including FAST++, FAST-CS, FAST-pw, and FAST-all. FAST++ and FAST-CS utilize k-means++ clustering (Arthur and Vassilvitskii, 2007) and constructed coresets (Bachem et al., 2018). In contrast, FAST-pw and FAST-all leverage minhashing and locality-sensitive hashing (Leskovec et al., 2014). FAST-R can be applied in two scenarios, namely the fixed-budget and adequate scenario. The fixed-budget version of FAST-R minimizes the test suite using a pre-defined minimization budget, while the adequate version prioritizes preserving the requirement coverage of the minimized test suite over adhering to the budget. One limitation of FAST-R is that it cannot simultaneously guarantee both minimization budget and full requirement coverage. We compare RTM against FAST-R under both scenarios where FAST-R is applicable (i.e., fixed-budget and adequate test suite minimization). For FAST-R, we relied on the publicly available replication package [^3] https://github.com/ICSE19-FAST-R/FAST-R provided by its authors to ensure a fair and consistent comparison.

### v2

@@ 3.4. Evolutionary Search for Test Suite Minimization
As discussed in Section 2.2, the TSM in our context is an NP-hard problem. Therefore, to efficiently find a near-optimal, feasible solution, we apply a GA–the most widely used evolutionary search algorithm in software testing and the approach adopted by the state-of-the-art (SOTA) TSM work (i.e., ATM (pan2023atm) and LTM (DBLP:journals/tse/PanGB24))–which has demonstrated high effectiveness and efficiency for TSM. To tailor GA to our problem, we define the individuals (minimized test suites), the initialization strategy, selection, crossover, and mutation operators, and the fitness function that guides the search. The details are given below.
@@ 3.4.1. Problem Formalization
Given the test suite before minimization $\mathcal{T}$, requirements set $\mathcal{R}$ with $r$ requirements, budget $M\in(0,1]$, the minimization problem is defined as a fixed-size subset selection with $100\%$ requirement coverage as the constraint. Each individual (i.e., subset $\mathcal{T}^{\prime}$) is represented as a binary vector:
@@
Given the minimization budget ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}\in(0,1]$, the number of test cases in the subset $\mathcal{T}^{\prime}$ is constrained by:
(9) $\sum_{i=1}^{m}x_{i}=\max\!\left(r,\operatorname{round}\!\left(m\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}\right)\right)$
where $\mathrm{round}()$ denotes rounding the number to the nearest integer. Note that the number of test cases in $\mathcal{T}^{\prime}$ should be higher than or equal to the total number of requirements.
@@
Given that the objective of the search process is to minimize the similarity values between test case pairs in the minimized test suite, the fitness function is formulated as:
(11) $\mathrm{Fitness}(\mathcal{T}^{\prime}_{n})=\frac{1}{n}\sum_{i=1}^{n}\biggl(\max_{\begin{subarray}{c}1\leq j\leq n\\ j\neq i\end{subarray}}\operatorname{NormSim}(t_{i},t_{j})\biggr)^{2},\quad\text{where }\mathcal{T}^{\prime}_{n}=\{t_{1},\dots,t_{n}\}.$
where $\mathcal{T}^{\prime}_{n}$ denotes the subset containing $n$ test cases, and $\operatorname{NormSim}(t_{i},t_{j})$ represents the normalized similarity score for the test case pair $(t_{i},t_{j})$. Individuals with lower fitness values are better, indicating more diverse test cases with lower similarity.
@@ 3.4.2. Initialization strategies
Our minimization problem is defined with two hard constraints: (1) the percentage of selected test cases in the subset should satisfy the minimization budget, and (2) the subset should preserve $100\%$ requirement coverage. To efficiently guide the search for valid subsets, we investigate three strategies to initialize a set of valid subsets for the GA to use as a base for further search. Note that, in our dataset, as is often the case, the relationship between test cases and requirements is many-to-one. Consequently, one experimental setting of our minimization problem is that the pre-defined number of selected test cases ($|\mathcal{T}^{\prime}|$) must not be smaller than the total number of requirements ($|\mathcal{R}|$ ), thus satisfying the full requirement coverage constraint when combined with the initialization strategy.
  • Strategy 1. Iterative Selection.
    • This strategy iteratively selects one test case randomly under each requirement, continuing this process until the percentage of selected test cases reaches the minimization budget.
  • Strategy 2. Initial Requirement + Random.
@@
    • Step 2: Randomly select the remaining test cases without considering requirements until the percentage of selected test cases reaches the minimization budget.
  • Strategy 3. Proportional Selection.
@@
    • Step 2: If the percentage of selected test cases across all requirements is less than or greater than the minimization budget, iteratively add or reduce the number of selected test cases per requirement, while ensuring every requirement is still covered by at least one test case, until the desired minimization budget is reached.
3.4.3. Search Process
After initializing a population with valid individuals (subsets), GA (golberg1989genetic) iteratively evolves this population using selection, crossover, and mutation operators towards better individuals with lower fitness values. In each generation, offspring are created by recombining and mutating the genes (i.e., selected test cases) of existing individuals. Each individual is evaluated by computing its fitness, and those with lower fitness are selected for the next generation. The evolutionary process is repeated until convergence, defined as a decrease in the fitness score of less than $0.0025$ across generations. The final output is the best subset identified during the search, which satisfies all constraints and minimizes the redundancy among selected test cases, indicated by the lowest fitness value.
Specifically, we adopt the following three genetic operators:
@@ 3.4.3. Search Process
  • Crossover: We used the customized crossover operator (pymoo) that ensures the offspring maintains a fixed subset size. It first includes all test cases shared by both parent subsets, then adds test cases from either parent subset until the minimization budget is met.
  • Mutation: We used an inversion mutation operator (pymoo) that randomly selects a segment of the subset and reverses its order. Since a subset is represented as a binary vector, this approach preserves the subset size.
@@
We follow the published guidelines  [^1] https://www.obitko.com/tutorials/genetic-algorithms/recommendations.php used in prior work (DBLP:journals/tse/PanGB24) for setting GA hyperparameters. Specifically, we use a population size of $100$, a mutation rate of $0.01$, and a crossover rate of $0.90$. For each generation, the size of each solution (i.e., subset) remains fixed at the minimization budget.
@@ 1. RQ1:
    This RQ aims to identify the optimal configuration of RTM that achieves the best performance in terms of FDR. Specifically, a RTM configuration entails the selection of a preprocessing method (Section 3.1), an embedding technique (Section 3.2), a similarity measure (Section 3.3), and an initialization strategy (Section 3.4.2).
@@ 2. RQ2:
    How does RTM compare to baseline TSM techniques?
    This RQ aims to compare the performance of RTM with two baseline approaches, namely Random Minimization and FAST-R (cruciani2019scalable), under various minimization budgets (i.e., 10%, 20%, …, 90%).
@@ 3. RQ3:
  4. RQ4:
    How does RTM scale with test suite size compared to baseline TSM techniques?
    This RQ aims to assess the scalability of RTM compared to the baseline techniques. Specifically, we investigate how the run time— including preparation time, search time and total minimization time—of RTM scale with test suite size, compared to baseline techniques.
@@ 4.2. Dataset
We evaluated RTM and compared it to baselines on a dataset collected from our industry partner, an automotive system company. There are $736$ test cases across seven test runs. As shown in Figure 1, the test cases consist of test steps that validate the behavior of the automotive system related to a specific DTC. Each test case covers one requirement, covering a total of $54$ requirements, and may detect multiple faults, detecting a total of $220$ unique faults.
Note that we identified the faults, which are the root causes of the failing test cases (e.g., parameter value mismatches, connection errors in specific components), using test execution logs under the guidance of the system engineers. For each test case, we first extracted failure messages from the test execution logs and then grouped those triggered by the same root cause into faults.
Test suites for different redundancy levels. In this dataset, a single fault can be detected by multiple test cases, indicating redundancy in the test suite. We define the redundancy level for the test suite as follows:
@@
For example, for a test suite $\mathcal{T}$ with three test cases that detect $3$, 22, and $3$ faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is $4$, the redundancy level of $\mathcal{T}$ is therefore $(3+2+3)/4=2$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
To assess the impact of the test suite redundancy level on the performance of RTM, we generate $10$ diverse test suites for each redundancy level using Integer Linear Programming (ILP)  [^2] https://pypi.org/project/PuLP/ and GA. This process starts by utilizing ILP to search for $100$ test suites with different test suite sizes while satisfying three constraints: (1) $100\%$ requirement coverage, (2) $100\%$ fault coverage, and (3) a specific test redundancy level. We set $15$ different redundancy levels, ranging from $4.5$ to $11.5$, with an interval of $0.5$. Note that no test suite satisfies a redundancy level below $4.5$ under constraints (1) and (2), and the highest redundancy level is $11.86$, which is observed in the original test suite. Then, to enhance the diversity of the dataset for experiments, for the $100$ test suites under each redundancy level, we employ GA to search for the most diverse $10$ test suites that minimize the sum of the set similarity between test suites, which is defined as the Jaccard Similarity between two test suites:
@@
where $|T_{i}$ $\cap$ $T_{j}|$ denotes the number of test cases in the intersection of test suite $T_{i}$ and $T_{j}$. $|T_{i}\cup T_{j}|$ denotes the number of test cases in the union of test suites $T_{i}$ and $T_{j}$.
@@ 4.3. Text Representations
Inspired by prior work on vector representations of requirements (zhao2025machine; abbas2025requirements) and test cases (viggiato2022identifying), we employ a variety of text embedding techniques to convert test cases into vector representations. Specifically, we use TF-IDF, USE, Amazon Titan Text Embedding V2, and LongT5 to generate sentence-level embeddings, while Word2Vec, GloVe, and FastText are used to obtain word-level embeddings. Due to limitations in computational resources, data privacy constraints, and restrictions on input token length, we were unable to experiment with models such as BERT (devlin2019bert) and Sentence-BERT (DBLP:conf/emnlp/ReimersG19).
@@
TF-IDF. TF-IDF has demonstrated promising performance in identifying similar test cases expressed in natural language (viggiato2022identifying). We also employed TF-IDF to extract numerical vector representations of the test cases. For each word, we calculated its importance within a single test case relative to its occurrence across all other test cases.
@@
USE (Universal Sentence Encoder) (DBLP:journals/corr/abs-1803-11175). USE encodes text into high-dimensional vectors, capturing semantic information at the sentence level, making it particularly effective for tasks such as semantic similarity, text classification, clustering, and information retrieval (DBLP:journals/corr/abs-1803-11175).
@@
LongT5 (DBLP:conf/naacl/GuoAUONSY22). LongT5 is an extension of the T5 architecture designed to efficiently handle long sequences (DBLP:conf/naacl/GuoAUONSY22). It incorporates sparse attention and reversible layers to reduce memory usage, making it well-suited for tasks like long-document summarization and retrieval-augmented generation. We selected the local attention mechanism for LongT5 because it yielded a higher FDR for TSM than the global attention mechanism.
@@
Amazon Titan Text Embedding V2 (aws2024titantv2). For LLM-based embeddings, we use Amazon Titan Text Embedding V2, provided by our industrial partners. The model takes test case steps as input and produces a $1,024$-dimensional vector representation for test cases.
@@
Word2Vec (DBLP:journals/corr/abs-1301-3781). Word2Vec has been shown to be effective for identifying similar test cases (viggiato2022identifying). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ for Word2Vec, as this configuration yielded the highest FDR for TSM. This was determined through experiments evaluating both the CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$, in increments of $2$. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
@@
GloVe (DBLP:conf/emnlp/PenningtonSM14). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (DBLP:conf/emnlp/PenningtonSM14). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of $2$ for GloVe based on its performance in hyperparameter tuning experiments with window sizes ranging from $2$ to $10$, in increments of $2$. As in Word2Vec, we used GloVe to generate a $300$-dimensional numerical vector representation for each word in the test cases.
@@
FastText (DBLP:journals/tacl/BojanowskiGJM17). FastText is a word embedding model proposed by Bojanowski et al. (DBLP:journals/tacl/BojanowskiGJM17), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of $8$ as the FastText configuration, based on experiments evaluating both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
@@ 4.4. Baselines
We compare RTM against (1) Random Minimization, a standard TSM baseline (pan2023atm), (2) FAST-R approaches (cruciani2019scalable), a set of novel TSM approaches with two versions: fixed-size budget version and adequate version. The former satisfies the pre-defined minimization budget constraint, whereas the latter satisfies the full requirement coverage constraint. While being a widely used approach for TSM (khan2018systematic), Greedy-based TSM approaches iteratively select the test cases that maximize the coverage until the full coverage constraint is satisfied (miranda2017scope; noemmer2019evaluation). However, in our dataset, each test case covers only one requirement, rendering this option equivalent to Random Minimization with a requirement coverage constraint, as described below.
@@ 4.4.1. Random Minimization
Since the goal of TSM is to reduce the number of test cases and, consequently, lower testing costs, the most straightforward approach is to randomly remove test cases to meet the minimization budget (DBLP:journals/tse/AliBHP10). We evaluate the performance of RTM against Random Minimization with and without the requirement coverage constraint, denoted as RM-Req and RM-NoReq, respectively. For the version with the constraint, we applied the same selection strategy as GA initialization described in Section 3.4.1. For the version without the constraint, we randomly select $k$ test cases from the test suite, ensuring that $k$ adheres to the minimization budget.
4.4.2. FAST-R (cruciani2019scalable)
FAST-R is a family of similarity-based TSM approaches, including FAST++, FAST-CS, FAST-pw, and FAST-all. FAST++ and FAST-CS utilize k-means++ clustering (DBLP:conf/soda/ArthurV07) and constructed coresets (DBLP:conf/kdd/BachemL018). In contrast, FAST-pw and FAST-all leverage minhashing and locality-sensitive hashing (DBLP:books/cu/LeskovecRU14). FAST-R can be applied in two scenarios: the fixed-budget and adequate scenarios. The fixed-budget version of FAST-R minimizes the test suite within a predefined budget, while the adequate version prioritizes preserving the requirement coverage of the minimized test suite over adhering to the budget. One limitation of FAST-R is that it cannot simultaneously guarantee both a minimization budget and full requirement coverage. We compare RTM against FAST-R under both scenarios where FAST-R is applicable (i.e., fixed-budget and adequate test suite minimization). For FAST-R, we relied on the publicly available replication package [^3] https://github.com/ICSE19-FAST-R/FAST-R provided by its authors to ensure a fair and consistent comparison.